OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). sin2 x - cos2 x = 0

OpenStudy (anonymous):

sry no lol

OpenStudy (anonymous):

and you should convince yourself why it is fine to divide by cos2x . ( it is fine. )

OpenStudy (anonymous):

wud it b pie/4, pie/6 ?

OpenStudy (anonymous):

wait im stupid

OpenStudy (anonymous):

tan(2x) = 1 !!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

so the general solution for tan(x) = tan(y) x = y + pi*k so in our case: tan(2x) = tan(pi/4) 2x = pi/4 + pi*k x = pi/8 + (pi/2) * k

OpenStudy (anonymous):

pie/4 just then

OpenStudy (anonymous):

no

OpenStudy (anonymous):

u got over 8?

OpenStudy (anonymous):

pi/8 , 5pi/8 , 9pi/8 , 13pi/8

OpenStudy (anonymous):

i first wrote tan(2x) = 0 which was WRONG

OpenStudy (anonymous):

pi/4, 3pi/4, 5pi/4, 7pi/4

OpenStudy (anonymous):

no, x=pi/8 , 5pi/8 , 9pi/8 , 13pi/8

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

wait lol was it :

OpenStudy (anonymous):

\[\sin(2x) - \cos(2x) = 0\]

OpenStudy (anonymous):

or \[\sin^2(x) - \cos^2(x) = 0\]

OpenStudy (anonymous):

cause i solved for the FIRST one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it said i got it right for the four i picked? weird

OpenStudy (anonymous):

but what was the correct question i wrote here two

OpenStudy (anonymous):

i think i solved for the wrong one

OpenStudy (anonymous):

im lost sry its ok bc i got it right so its all good :)

OpenStudy (anonymous):

cause if the question was : sin^2(x) - cos^2(x) = 0 then we get tan^2(x) = 1 tan(x) = +/-1 so 1. tan(x) = 1 tan(x) = tan(pi/4) x = pi/4 + pi * k x = pi/4 , 5pi/4 2.tan(x) = -1 tan(x) = tan(-pi/4) x = -pi/4 + pi*k x = 3pi/4 , 7pi/4 so : x = pi/4, 3pi/4, 5pi/4, 7pi/4

OpenStudy (anonymous):

and i thought it was sin(2x) - cos(2x) = 0